Solve the system by the addition or substitution method.
x - 6y = -46
-3x - 7y = -62

Answer by sofiyac(695)

x - 6y = -46
-3x - 7y = -62 multiply first one by 3 and add to second one
***
3x - 18y = -138
-3x - 7y = -62 add them
************
-25y=-200 divide each side by -25
y=8
plug into either of the equations
x-6*8=-46
x-48=-46
x=2
(2,8)